OpenStudy (anonymous):

Solve using transformations. 13 + m = -96 The answer is -109, is that correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great, can you help me with the other questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thanks, the next question is Solve using transformations. n - (-16) = 75 n = 59

OpenStudy (anonymous):

59 + 16 = 75 right

OpenStudy (anonymous):

your answer was right...its 59

OpenStudy (anonymous):

60+ 16 = 76 76 -1 = 75 yes

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

59 is right

OpenStudy (anonymous):

Thanks, and no prob alex, you're helping me out :)

OpenStudy (anonymous):

The next question is Solve using transformations. | f | - 7 = 15 | f | = 22 or -22